25) Calculate the Internal Rate of Return based on the following data.

       Original Investment: $1,500

year net cash flow NPV discount 10% factor discounted cash flow @10% calculate NPV discount @30% facotr discounted cash flow @30% calculate
1 $2000 .8900 .3300
2 $2000 .7800 .2200
3 $1000 .6800 .0600
Answer #1

ANSWER

Year

Net Cash Flow ($)

NPV Discount Factor @10 %

Discounted Cash Flow @ 10%

NPV Discount Factor @30 %

Discounted Cash Flow @ 30%

A

B

C

D = B x C

E

F = B x E

1

2000

0.8900

1780

0.3300

660

2

2000

0.7800

1560

0.2200

440

3

1000

0.6800

680

0.0600

60

Total

4020

1160

Internal Rate of Return or Implicit Rate of Return where Present Value of cash inflow during the given 3 years = $ 1,500

Using Interpolation method, we get

X – X1 / X2 – X1 = Y – Y1 / Y2 – Y1

X – 10 / 30 – 10 = 1500 – 4020 / 1160 – 4020

Using BODMAS we get,

X – 10/20 = -2520/-2860

X = (0.88 x 20) + 10

X = 27.62 %

Where,

X = Desired rate of IRR

X1 = 1st assumed rate

X2 = 2nd assumed rate

Y1 = Cumulative PV when, Discount factor rate is X1

Y2 = Cumulative PV when, Discount factor rate is X2

Note: - The rates given in question as

Rate

Year 2

Year 2

Year 3

10%

0.8900

0.7800

0.6800

30%

0.3300

0.2200

0.0600

Have been duly accounted in question without first authenticating the same based on applicable rate percentage.